# Context-Aware Code Wiring Recommendation with LLM-based Agent

### v1

@@ IV-F RQ1: Outperform the SOTA
TABLE II: Comparison with Baselines
#EM #Rec $EM_{Precision}$ $EM_{Recall}$
IDEA 79 189 41.8% 35.7%
ExampleStack 4 10 40.0% 1.8%
GPT-4o-mini 101 145 69.7% 45.7%
Qwen2.5-Coder-14B 110 157 70.1% 49.8%
Qwen2.5-Coder-32B 149 197 75.6% 67.4%
Qwen-max 124 159 78.0% 56.1%
DeepSeek-V3 129 171 75.4% 58.4%
WIRL - Locator 110 157 70.1% 49.8%
WIRL - Collector 169 190 88.9% 76.5%
WIRL - Completer 146 189 77.2% 66.1%
WIRL 199 217 91.7% 90.0%
In Table II, the first column lists the evaluated approaches. The second and third columns indicate, respectively, the number of cases in which the recommended code elements exactly match the ground truth and the number of cases in which a recommendation is made by the approach. The final two columns, $EM_{Precision}$ and $EM_{Recall}$, report how often the recommendations are correct and how many of the correct substitutions are successfully identified, respectively (see Section IV-D for detailed definitions). From Table II, we make the following observations:
@@
  • Compared to IDEA, WIRL achieves improvements of 54.3 percentage points in $EM_{Recall}$ and 49.9 percentage points in $EM_{Precision}$. Against ExampleStack, it shows even greater gains: 88.2 and 51.7 percentage points in $EM_{Recall}$ and $EM_{Precision}$, respectively. When benchmarked against the advanced LLM model Qwen2.5-Coder-32B, WIRL still delivers notable improvements of 22.6 and 13.7 percentage points, further demonstrating its superior effectiveness.
@@
  • WIRL also substantially improves upon the performance of directly using a raw LLM (Qwen2.5-Coder-14B). Without the academic design described in Section III, the baseline LLM achieves only 49.8% $EM_{Recall}$ and 70.1% $EM_{Precision}$. With the integration of the external toolkit and structured prompt design, WIRL improves $EM_{Recall}$ and $EM_{Precision}$ by 40.2 and 21.6 percentage points, respectively, highlighting the value of its architectural innovations.
  • The three types of tools in WIRL contribute differently to its performance: locator tools have the greatest impact, followed by completer tools, with collector tools contributing the least. Locator tools contribute the most because the collector and completer tools rely on the location information provided by them. Removing the locator tools will reduce the system to relying solely on raw LLMs.
We further analyzed the reasons behind WIRL’s superior performance compared to baseline approaches, with a particular focus on IDEA. During the evaluation of IDEA, we recorded not only the top-ranked recommendation but also the complete list of suggestions. Our quantitative and qualitative analysis revealed that in 84.1% of the cases (159 out of 189), IDEA did include the correct recommendation, but it was not ranked first. Statistically, the correct answer appeared in the top position in only 49.7% of those 159 cases. Notably, in some instances, the recommendation list contained more than 100 items, requiring developers to scroll extensively before locating the correct option. These findings indicate that while IDEA’s underlying recommendation algorithm, originally designed for renaming suggestions, can potentially address the code wiring problem, it lacks an effective ranking mechanism to prioritize relevant suggestions. This inefficiency increases the cognitive load and selection time for developers. In contrast, WIRL generates a single high-confidence recommendation for each unresolved element, achieving a high $EM_{Precision}$ of 91.7%. This direct substitution approach not only reduces the time and effort required from developers but also allows them to focus more on adapting business logic rather than sifting through lengthy suggestion lists. Furthermore, IDEA failed to generate any recommendation in 14.5% of the cases (32 out of 221). These failures occurred when the unresolved elements were expressions rather than variables or literals, i.e., cases that exceed the intended scope of IDEA’s renaming functionality. In contrast, WIRL successfully handled such cases through its use of static code analysis, showcasing its broader applicability and robustness.
@@
For ExampleStack, we conducted a manual analysis of the input data and identified the key limitation behind its unsatisfied performance. The core rationale of ExampleStack depends on a pre-collected database of historical modifications of code snippets from Stack Overflow to GitHub. Consequently, its effectiveness is heavily constrained by the presence or absence of identical or similar historical modifications in the database. This reliance significantly hampers its applicability to previously unseen cases, which are common in real-world development scenarios. In fact, within the $CWEvaluation$, only 6 code wiring instances involving 18 unresolved elements had similar modifications present in the original database provided by ExampleStack [12]. Across all test cases, ExampleStack made 10 recommendations, only 4 of which were correct. In contrast, WIRL leverages the reasoning and generation capabilities of large language models (LLMs) to recommend context elements dynamically and from scratch. This design enables WIRL to operate independently of historical reuse data, thereby overcoming the fundamental limitation of ExampleStack and improving its generalization and effectiveness across diverse code wiring scenarios.
Regarding the performance of raw LLMs, their relatively low $EM_{Recall}$ can be attributed to two main factors. First, the length of the input class significantly impacts the ability of LLMs to make accurate recommendations. When the class size exceeds the model’s maximum token limit, the LLM cannot process the full context, resulting in failure to generate any meaningful output. For instance, the maximum token limit for DeepSeek-V3 is 64K tokens. If a class exceeds this size, DeepSeek-V3 becomes incapable of handling the instance entirely. Second, raw LLMs sometimes fail to identify all unresolved elements within the code. As illustrated in Fig. 3, there are two unresolved elements, i.e., “list” and “target”, with their corresponding context elements being “listView” and “mCommentListPosition”. While Qwen2.5-Coder-14B successfully identified “list” and recommended “listView”, it failed to detect “target”, thus missing a key aspect of the task. In contrast, WIRL, by incorporating compiler information, accurately identified both unresolved elements (including all six references to “list”) and successfully recommended the appropriate context elements, “listView” and “mCommentListPosition”, by leveraging the reasoning capability of LLMs.
@@
The relatively low $EM_{Precision}$ of raw LLMs can be attributed to their difficulty in analyzing long and complex contexts to correctly identify mapping relationships, particularly when the appropriate context element is not explicitly present in the immediate code. As illustrated in Fig. 4, the unresolved element is encoding, and its corresponding context element is “Charset.defaultCharset()”. It is challenging for raw LLMs to recommend “Charset.defaultCharset()” without the context information that the “Charset” class provides a static method “defaultCharset()” that returns an instance of “Charset”. In contrast, WIRL addresses this limitation by leveraging its toolkit. The initial analysis of the context (is_argument tool) determines the type of encoding, which is “Charset”. Then, it uses the get_method_names tool to retrieve methods in the “Charset” class that return the same type. This process successfully identifies “defaultCharset()”, enabling WIRL to include it in the prompt for the final completion module, which then makes the correct recommendation. Additionally, raw LLMs occasionally generate substitutions beyond the unresolved elements, introducing unexpected recommendations and increasing the number of false positives, which further impacts $EM_{Precision}$.
@@ IV-G RQ2: Time Efficiency
TABLE III: Time, Token Consumption, and Monetary Cost
Avg. Time Cost (ms) Avg. #Input Tokens Avg. #Output Tokens Avg. #Total Tokens Avg. Monetary Cost (CNY)
IDEA 3.7 - - - -
ExampleStack 729.1 - - - -
GPT-4o-mini 3,930.4 2,769.4 27.4 2,796.8 0.0031
Qwen2.5-Coder-14B 4,624.5 2,769.4 28.3 2,797.7 0.0057
Qwen2.5-Coder-32B 6,799.4 2,769.4 43.8 2,813.2 0.0058
Qwen-max 5,393.6 2,769.4 24.9 2,794.3 0.0069
DeepSeek-V3 11,489.9 2,769.4 22.0 2,791.4 0.0057
WIRL 5,255.6 3,006.7 108.5 3,115.2 0.0067
In this research question, we examine the time efficiency of WIRL compared to other baseline approaches on the $CWEvaluation$. The comparison results are summarized in Table III, where the second column reports the average execution time per code wiring instance. It is worth noting that the reported time costs are averaged over five independent runs to ensure robustness and consistency. From Table III, we make the following observations:
@@
To gain deeper insights into the time cost distribution of WIRL and the baseline approaches, we visualized the results using a violin plot, as shown in Fig. 5. The plot reveals that WIRL is capable of handling the majority of code wiring instances in under 5 seconds. Its median time cost is 4,201.5ms, which is comparable to that of Qwen2.5-Coder-14B (3,977.5ms), and substantially lower than that of DeepSeek-V3 (7,983.5ms). This demonstrates WIRL’s ability to deliver high-quality recommendations with competitive latency.
@@ IV-H RQ3: Token Consumption and Monetary Cost
In this research question, we evaluate the token usage and monetary cost of WIRL compared to other LLM-based baselines on the $CWEvaluation$. The comparison results are summarized in Table III. Columns three to five report the average number of input tokens, output tokens, and total tokens per code wiring instance, respectively, while the final column shows the corresponding average monetary cost. All figures are averaged over five independent experimental runs.
As shown in Table III, WIRL consumes slightly more tokens than Qwen2.5-Coder-14B, with an average increase of 237.3 input tokens and 80.2 output tokens. This results in a marginally higher monetary cost, i.e., approximately 0.001 CNY more per instance. Surprisingly, although GPT-4o-mini incurs the lowest monetary cost, its performance is relatively poor, as shown in Table II. The increase in token usage is attributed to WIRL’s iterative design, where it progressively collects and integrates contextual information into the prompt before generating a final recommendation. As the output from one iteration is passed as input to the next, the cumulative token count naturally exceeds that of the single-pass LLM baselines.
@@ V Discussion
V-A Threats to Validity
A potential threat to external validity lies in the limited size and diversity of the dataset used for evaluation. To mitigate this risk, we leveraged a publicly available dataset and performed a thorough manual inspection to ensure its quality. Additionally, we have open-sourced both the dataset and the evaluation results [8] to facilitate replication and further validation by the research community.

### v2

@@ IV-F RQ1: Outperform the SOTA
TABLE III: Comparison against Baselines
Baselines #EM #Rec $EM_{P}$ $EM_{R}$ Latency
IDEA 79 189 41.8% 35.7% 3.7ms
ExampleStack 4 10 40.0% 1.8% 729.1ms
CoEdPilot 21 32 65.6% 9.5% 2,417.4ms
GrACE 163 184 88.6% 73.8% 2,873.4ms
GPT-4o-mini 101 145 69.7% 45.7% 3,930.4ms
Qwen2.5-Coder-14B 110 157 70.1% 49.8% 4,624.5ms
Qwen2.5-Coder-32B 149 197 75.6% 67.4% 6,799.4ms
Qwen-max 124 159 78.0% 56.1% 5,393.6ms
DeepSeek-V3 129 171 75.4% 58.4% 11,489.9ms
WIRL 199 217 91.7% 90.0% 5,255.6ms
In Table III, the first column lists the evaluated approaches. The second and third columns indicate, respectively, the number of cases in which the recommended code elements exactly match the ground truth and the number of cases in which a recommendation is made by the approach. The fourth and fifth columns, $EM_{Precision}$ and $EM_{Recall}$, report how often the recommendations are correct and how many of the correct substitutions are successfully identified, respectively (see Section IV-D for detailed definitions). From Table III, we make the following observations:
@@
  • Compared to IDEA, WIRL achieves improvements of 54.3 percentage points in $EM_{Recall}$ and 49.9 percentage points in $EM_{Precision}$. Against ExampleStack, it shows even greater gains: 88.2 and 51.7 percentage points in $EM_{Recall}$ and $EM_{Precision}$, respectively. Compared to advanced code editing approaches, i.e., GrACE and CoEdPilot, WIRL achieves improvements of 16.2 and 80.5 percentage points in $EM_{Recall}$ and 3.1 and 26.1 percentage points in $EM_{Precision}$, respectively. When benchmarked against the advanced LLM model Qwen2.5-Coder-32B, WIRL still delivers notable improvements of 22.6 and 13.7 percentage points, further demonstrating its superior effectiveness.
@@
  • WIRL also substantially improves upon the performance of directly using a raw LLM (Qwen2.5-Coder-14B). Without the academic design described in Section III, the baseline LLM achieves only 49.8% $EM_{Recall}$ and 70.1% $EM_{Precision}$. With the integration of the external toolkit and structured prompt design, WIRL improves $EM_{Recall}$ and $EM_{Precision}$ by 40.2 and 21.6 percentage points, highlighting the value of its architectural innovations.
IV-F1 Analysis Over IDEA
First, IDEA’s suggestion mechanism, designed for renaming, lacks effective ranking mechanism for code wiring. Our analysis revealed that while IDEA’s suggestion list contained the correct substitution in 84.1% of cases (159 out of 189), it was often buried among a large number of irrelevant options, sometimes over 100, requiring significant developer effort to locate. Only 49.7% candidates in the top position are correct answers. This increases cognitive load and slows down the adaptation process. In contrast, WIRL provides a single, high-confidence recommendation for each unresolved element with 91.7% precision, turning a tedious search into a simple validation step. Second, IDEA failed to generate any recommendations in 14.5% of cases. These failures typically occurred when the unresolved elements are involved with complex expressions rather than simple variable s, a scenario that falls outside the scope of its renaming functionality. WIRL, by leveraging static code analysis, successfully handles these more complex cases, demonstrating its broader applicability and robustness.
IV-F2 Analysis Over ExampleStack
@@
The core limitation of ExampleStack is its dependency on a pre-collected database of historical code modifications. Its effectiveness is therefore constrained to scenarios that have been seen before, which significantly hampers its performance on novel problems common in real-world development. In our evaluation, only 6 code wiring instances within $CWEvaluation$ had similar historical modifications in ExampleStack’s database [6]. Across all test cases, ExampleStack attempted 10 recommendations, with only 4 being correct. In contrast, WIRL leverages the reasoning capabilities of LLMs to generate solutions dynamically and from scratch for each specific context. This design allows WIRL to operate independently of historical data, giving it far greater generalization and effectiveness across diverse and unseen code wiring scenarios.
IV-F3 Analysis Over Code Editing Methods
As presented in Table III, WIRL surpasses the performance of advanced code editing methods like GrACE and CoEdPilot. This superiority can be attributed to two primary factors: First, these code editing methods are fundamentally designed to leverage a rich history of prior code edits. The code wiring scenario, however, involves pasting code from external sources, a context that inherently lacks such a historical record. This foundational mismatch significantly curtails the effectiveness of code editing methods. Second, the core methodologies of these tools are not well-suited for code wiring. CoEdPilot’s key modules for analyzing prior and subsequent edits are ineffective in this context, as there is little relevant edit history to analyze. Consequently, its generation module, which is fine-tuned specifically for edit-history-rich tasks, is ill-equipped for the distinct challenge of code wiring. GrACE, utilizing GPT-4o in a zero-shot setting, achieves impressive results by combining the powerful completion capabilities of LLMs with a carefully engineered prompt. This design philosophy is highly analogous to that of WIRL, and its success further validates the effectiveness of our approach.
IV-F4 Analysis Over Raw LLMs
WIRL’s superiority over raw LLMs is evident in both recall and precision. The lower $EM_{Recall}$ of raw LLMs can be attributed to two main factors. First, their fixed context windows prevent them from processing large classes that exceed token limits, resulting in complete failure. Second, they often fail to identify all unresolved elements within the code. As illustrated in Fig. 3, Qwen2.5-Coder-14B successfully identified “list” and recommended “listView”, but it failed to detect “target”. In contrast, WIRL uses compiler information to reliably identify all unresolved elements, overcoming these limitations.
@@ IV-F4 Analysis Over Raw LLMs
The reduced $EM_{Precision}$ of raw LLMs stems from their difficulty in analyzing complex contexts, especially when the substitution is not a simple variable but requires synthesizing code, e.g., a method call. As shown in Fig. 4, a raw LLM is unlikely to deduce the correct substitution “Charset.defaultCharset()” because it lacks awareness of the class’s available static methods. In contrast, WIRL addresses this limitation by leveraging its toolkit to first identify the required type (e.g., “Charset”) and then retrieve relevant methods (e.g., “defaultCharset()”), supplying the LLM with the necessary information to make a correct recommendation. Additionally, raw LLMs sometimes generate extraneous substitutions, increasing false positives and further lowering precision.
IV-F5 Failure Analysis for WIRL
While WIRL demonstrates strong performance, analyzing its failures reveals its current limitations. We use a representative case testing with different base models (Fig. 5) to illustrate this boundary. In this case, WIRL must resolve three elements: “url1”, “user”, and “password”. It successfully maps the simpler elements, “user” and “password”, to their corresponding context variables, “properties.getUser()” and “properties.getPassword()”. This highlights its proficiency in handling direct mappings. The failure occurred with the element “url1”. Its correct substitution is not a simple variable but a deeply nested chain of method invocations. Constructing this requires advanced, multi-step compositional reasoning, a known frontier challenge for current LLMs and agent systems. The model needs to synthesize a complex code structure rather than just perform a direct mapping. This failure across various base models indicates that the limitation lies not only in the LLM’s capabilities but also in the current agent-tool interaction paradigm, which is less effective for tasks requiring the synthesis of complex, multi-step code from scratch.
@@ IV-G RQ2: Time Efficiency
In this research question, we examine the time efficiency of WIRL compared to other baseline approaches on the $CWEvaluation$. The comparison results are summarized in Table III, where the last column reports the average execution time per code wiring instance. It is worth noting that the reported time costs of raw LLMs are averaged over five independent runs to ensure robustness and consistency. From Table III, we make the following observations:
@@
  • Code editing methods, such as GrACE and CoEdPilot, exhibit moderate time efficiency, placing them between the fast, heuristic-based methods and the slower, raw LLM baselines. This performance is attributable to their reliance on pretrained language models, which are more computationally intensive than simple heuristics.
To gain deeper insights into the time cost distribution of WIRL and the raw LLMs, we visualized the results using a violin plot, as shown in Fig. 6. The plot reveals that WIRL is capable of handling the majority of code wiring instances in under 5 seconds. Its median time cost is 4,201.5ms, which is comparable to that of Qwen2.5-Coder-14B (3,977.5ms), and substantially lower than that of DeepSeek-V3 (7,983.5ms). This demonstrates WIRL’s ability to deliver high-quality recommendations with competitive latency.
@@ IV-H RQ3: Token Consumption and Monetary Cost
TABLE IV: Token Consumption, and Monetary Cost
LLMs Avg. #Input Tokens Avg. #Output Tokens Avg. #Total Tokens Avg. Monetary Cost (USD)
GPT-4o-mini 2,769.4 27.4 2,796.8 0.0004
Qwen2.5-Coder-14B 2,769.4 28.3 2,797.7 0.0008
Qwen2.5-Coder-32B 2,769.4 43.8 2,813.2 0.0008
Qwen-max 2,769.4 24.9 2,794.3 0.0010
DeepSeek-V3 2,769.4 22.0 2,791.4 0.0008
WIRL 3,006.7 108.5 3,115.2 0.0009
In this research question, we evaluate the token usage and monetary cost of WIRL compared to other LLM-based baselines on the $CWEvaluation$. The comparison results are summarized in Table IV. Columns 2-4 report the average number of input tokens, output tokens, and total tokens per code wiring instance, respectively, while the last column shows the corresponding average monetary cost. All figures are averaged over five independent experimental runs.
As shown in Table IV, WIRL consumes slightly more tokens than Qwen2.5-Coder-14B, with an average increase of 237.3 input tokens and 80.2 output tokens. This results in a marginally higher monetary cost, i.e., approximately 0.0001 USD more per instance. Surprisingly, although GPT-4o-mini incurs the lowest monetary cost, its performance is relatively poor (refer Table III). The increase in token usage is attributed to WIRL’s iterative design, where it progressively collects and integrates contextual information into the prompt before generating a final recommendation. As the output from one iteration is passed as input to the next, the cumulative token count naturally exceeds that of the single-pass raw LLMs.
@@
IV-I RQ4: Ablation Study
TABLE V: Ablation Study
Variants #EM #Rec $EM_{P}$ $EM_{R}$ Latency
WIRL 199 217 91.7% 90.0% 5,255.6 ms
w/o Locator 110 157 70.1% 49.8% 4,624.5 ms
w/o Collector 169 190 88.9% 76.5% 5,048.2 ms
w/o Completer 146 189 77.2% 66.1% 4,827.1 ms
w/o State Machine 198 217 91.2% 89.6% 7,297.1 ms
w/o Filtering Tools 179 205 87.3% 81.0% 5,679.7 ms
w/o Retrieval Tools 189 216 87.5% 85.5% 5,820.3 ms
To investigate the impact of each design choice on overall performance, we conducted an ablation study by creating several variants of WIRL. We first individually removed the three types of customized tools, Locator, Collector, and Completer, to assess their respective contributions (Rows 3-5 in Table V). Next, we evaluated the impact of key architectural components. We replaced the state machine with a fully agent-guided exploration strategy and individually removed the filtering tools (i.e., reserve_type_compatible_ones and sort_by_literal_similarity), and the retrieval tools (retrieve_identical_function_call and get_method_names). From Table V, we make the following observations:
  • Tool Contributions: Three tool types contribute differently to WIRL’s performance. The locator tools are the most critical, followed by the completer and then the collector tools. The paramount importance of the locator tools is because both the collector and completer depend on the precise location information they provide. Without locator, the system effectively degrades to a raw LLM.
  • Architectural Necessity: All architectural components are integral to the overall performance. The data clearly shows that removing any component, the state machine, filtering tools, or retrieval tools, results in a performance degradation. This highlights their crucial roles in providing valid contextual information to the agent. While the time efficiency remains stable across most variants, the removal of the state machine presents a notable exception.
  • State Machine Efficiency: The moderate latency increase observed when removing the state machine is a direct result of a proactive design choice to prevent aimless exploration, a common pitfall in agentic systems. We constrained the maximum iteration number to two, which limits the observable latency penalty in this study. Without this constraint, removing the state machine would lead to prolonged, multi-step exploration cycles and a significantly more substantial, and likely unacceptable, increase in latency.
The ablation study results confirm that each component is integral to WIRL’s performance and the architecture designs are also the key contributors to the efficiency of WIRL.
@@ V Discussion
V-A Placing WIRL in the Broader Context of Code Adaptation
Despite its specific focus, code wiring represents a frequent, critical, and foundational aspect of code adaptation. This claim is supported by two key factors:
  • Prevalence in Adaptations: The empirical study by Zhang et al. [3] on real-world code adaptations provides strong evidence for this. Their study revealed that a remarkable 88.0% of code reuses require adaptation. Through our further analysis, over one quarter (27.3%) of those adaptations involve code wiring operation, underscoring its high prevalence. Additionally, the study highlights that variable references and method invocations are central to the adaptation process, with 77.2% of statement-level code adaptations involving method invocations or variable references.
  • Task Complexity: While direct variable-to-variable mapping is relatively straightforward, the core challenge lies in mapping a variable to a complex expression, such as a deeply nested method invocation. This elevates the task from simple pattern matching to one requiring deeper program semantic understanding. A complex example is presented in Fig. 5 and the analysis is presented in Section IV-F5.
V-B Applicability Analysis
We evaluate WIRL’s real-world applicability across three key dimensions:
  • Accuracy: With a precision of 91.7%, WIRL significantly reduces the developer’s cognitive load by turning a tedious and error-prone manual adaptation process into a simple inspection and approval task in nine out of ten cases. This accuracy substantially outperforms the 71.8% of the previous state-of-the-art approach [9] in code snippet adaptation.
  • Latency: WIRL introduces a minimal computational overhead of only 0.6 seconds on top of the base model’s inference time. The average 5.25-second recommendation time, inflated by our experimental setup using online API calls, would be significantly reduced in a production environment with a locally deployed model or a dedicated API, making it well-suited for practical development workflows. To support self-hosted models or private APIs, the implementation of WIRL [30] allows for the configuration of a custom endpoint URL via the IDEA settings panel at File - Settings - Tools - WIRL AI Settings.
  • Cost: Economically, WIRL is highly efficient, with an average cost of just \$0.0009 per recommendation based on public API pricing. This negligible cost yields a high return on investment by saving valuable developer time.
V-C Dataset Representativeness Analysis
Although our dataset, $CWEvaluation$, is modest in size, its quality and representativeness are justified by:
  • Ecological Validity: $CWEvaluation$ is derived from genuine code reuse instances from the SOTorrent database filtered and validated by Zhang et al. [3], ensuring it reflects authentic development practices.
  • Project and Domain Diversity: It includes 29 distinct open-source projects from a wide range of domains, e.g., foundational libraries and popular mobile applications, enhancing the generalizability of our findings.
  • Structural Diversity: The code snippets’ structural properties, e.g., LOC, are comparable to those in larger datasets [3], and a distributional analysis can be found in our online website [31]. Besides, $CWEvaluation$ covers both file-level and project-level adaptation scenarios, ensuring its diversity and complexity.
V-D Versatility Analysis
WIRL is designed to be language-agnostic. Its core LLM-based reasoning and agentic control flow are language-independent. To adapt WIRL to a new language, only two language-specific components in its toolkit require modification: the Locator, which uses compiler information to find unresolved variables, and the Collector, which leverages Abstract Syntax Tree (AST) analysis to gather contextual information. This modular design ensures high generalizability.
V-E Threats to Validity